A car rental co. charges a daily fee plus a mileage fee. Joe was charged $429 for 5 days and 300 km., while Mary was charged 315 for 3 days and 500 km. What does the car rental co. charge per day and per kilometre ? Thanx for your help.
Answer by checkley77(12844)

5D+300M=429 MULTIPLY BY 3.
3D+500M=315 MULTIPLY BY -5 & ADD.
15D+900M=1,287
-15D-2500=-1,575
-------------------
-1600M=-288
M=-288/-1,600
M=.18 PRICE PER MILE DRIVEN.
5D+300*.18=429
5D+54=429
5D=429-54
5D=375
D=375/5
D=75 IS THE DAILY RENTAL RATE.
PROOF:
3*75+500*18=315
225+90=315
315=315